A G coil with its internal resistance 15 Ω, shows its full-scale deflection at 2 mA. How would you 

convert it into an A of range 0-5 A and a V of range 0-15 V? Give detailed calculations and draw 

the circuit diagrams for these conversions by mentioning the evaluated values of resistances.


1
Expert's answer
2021-07-06T18:09:04-0400

Resistance on galvanometer, "G=15\Omega"

current passing through the galvanometer, "I_g=2mA=2\times 10^{-3}A"

current across the ammeter, "I=5A"

use the formula, "S=\frac{I_g.G}{I-I_g}=\frac{2\times 10^{-3}\times 15}{5-0.002}=6m\Omega"

A short resistance of "6m\Omega" has to be connected in parallel to the galvanometer to convert it to an ammeter.